# Sensing-Enhanced Channel Estimation for Near-Field XL-MIMO Systems

### v1

@@ Abstract
Future sixth-generation (6G) systems are expected to leverage extremely large-scale multiple-input multiple-output (XL-MIMO) technology, which significantly expands the range of the near-field region. The spherical wavefront characteristics in the near field introduce additional degrees of freedom (DoFs), namely distance and angle, into the channel model, which leads to unique challenges in channel estimation (CE). In this paper, we propose a new sensing-enhanced uplink CE scheme for near-field XL-MIMO, which notably reduces the required quantity of baseband samples and the dictionary size. In particular, we first propose a sensing method that can be accomplished in a single time slot. It employs power sensors embedded within the antenna elements to measure the received power pattern rather than baseband samples. A time inversion algorithm is then proposed to precisely estimate the locations of users and scatterers, which offers a substantially lower computational complexity. Based on the estimated locations from sensing, a novel dictionary is then proposed by considering the eigen-problem based on the near-field transmission model, which facilitates efficient near-field CE with less baseband sampling and a more lightweight dictionary. Moreover, we derive the general form of the eigenvectors associated with the near-field channel matrix, revealing their noteworthy connection to the discrete prolate spheroidal sequence (DPSS). Simulation results unveil that the proposed time inversion algorithm achieves accurate localization with power measurements only, and remarkably outperforms various widely-adopted algorithms in terms of computational complexity. Furthermore, the proposed eigen-dictionary considerably improves the accuracy in CE with a compact dictionary size and a drastic reduction in baseband samples by up to 77%.
@@ I-A Related Works
Notably, an ideal dictionary needs to conform to the spatial characteristics of the channel in both distance and angle domains. This stipulation implies that acquiring localization information for both user equipments (UEs) and scatterers is a fundamental task in dictionary design for near-field CE [18]. Specifically, when prior localization information pertaining to the propagation environment is incorporated into the dictionary, it can significantly reduce the need for further pilots and subsequent baseband samples for near-field CE. In recent years, integrated sensing and communication (ISAC) has made significant progress, posing sensing as a promising functionality in future communication systems [19]. Thanks to the large aperture provided by XL-MIMO arrays, efficient location estimation [20] as well as movement tracking [21] of UEs and scatterers become possible. This motivates us to customize a novel sensing-enhanced CE scheme for near-field XL-MIMO.
There have been some initial attempts on localization via XL-MIMO or sensing-assisted CE. Near-field radar sensing utilizing orthogonal frequency division multiplexing (OFDM) waveforms was validated in [20], while [21] proposed a near-field tracking method by measuring the curvature of arrival of the spherical wavefront from the UE. Besides, a localization method based on the multiple signal classification (MUSIC) algorithm has also been proposed to decouple the estimation of angles of arrival (AoAs) and distances of UEs located in the near field [22]. However, an excessive number of additional baseband samples are required for localization when directly integrating such methods into existing CE schemes, which counteracts the reduction of samples for CE brought by localization. To further improve the utilization of pilot signals, a sensing-assisted CE scheme was then proposed in [23]. While the pilot signals are reused for both localization and CE procedures and therefore no baseband sampling is required for localization, fully-digital XL-MIMO arrays are essentially the prerequisite of existing sensing-assisted CE. This array architecture requires a large number of baseband samples (typically dozens of times the number of antennas on the XL-MIMO array) via power-hungry analog-to-digital converters (ADCs) and, hence, is far from practical. Furthermore, the computational complexity of classical localization algorithms, e.g., MUSIC, applied in [22] and [23] is exceptionally high, which is caused by massive matrix multiplications for super-resolution grid search on both distance and angle domains. Therefore, pragmatic CE schemes enhanced by localization with smaller sampling overhead and lower computational complexity still require further exploration.
@@ II System Model
In this section, we firstly describe the considered near-field transmission scenario and present the near-field channel model. Furthermore, the uplink sensing-enhanced CE protocol is illustrated. The power sensor-based and hybrid antenna array architectures are then introduced, whereas the corresponding uplink sensing and training models are further provided.
@@ II-A Channel Model
Consider a scenario with a UE array [^1] The proposed design in this paper can be extended to multi-user scenarios by assigning orthogonal pilots for different UEs. and a base station (BS) equipped with an XL-MIMO array operating in the TDD mode within its near-field region, where both the UE and BS are equipped with uniform linear arrays (ULAs) [^2] We consider ULA here for brevity, while it can be generalized to other antenna geometries. For example, it can be extended to uniform planar array (UPA) by applying Kronecker products to steering vectors in (2)., as shown in Fig. 1(a). The Cartesian coordinates of the $m$-th element of the UE’s antenna array and the $n$-th element of the BS antenna array are denoted by ${\bf r}_{\rm UE}^{(m)}=(x_{\rm UE}^{(m)},y_{\rm UE}^{(m)})^{T}$ and ${\bf r}_{\rm BS}^{(n)}=(x_{\rm BS}^{(n)},0)^{T}$, respectively. For uni-polarized antenna elements, the spherical wave spatial impulse response in the scalar form [25] is given by
@@ II-B Sensing-Enhanced CE Protocol and Power Senor-based XL-MIMO
We consider the uplink CE at the BS, which can be used to facilitate the downlink transmission. Conventionally, the UE transmits $\tau$ uplink pilot signals during the training stage, while the BS estimates the channel based on the received pilots [27, 13]. However, the pilot length $\tau$ and the resulting number of baseband samples at the BS can be exceedingly large due to limited RF chains in the hybrid array architecture and the complicated near-field channel model [18, 28].
In this paper, we propose a new sensing-enhanced CE protocol, where a single sensing time slot is inserted before the conventional pilot signal transmission, as shown in Fig. 2. In particular, in this additional slot, the UE transmits an uplink sensing signal to realize the localization of both the UE and scatterers. The main benefit of adding such a sensing time slot is that the estimated location coordinates can provide accurate prior localization information for facilitating efficient near-field CE. This not only reduces the pilot length $\tau$ in the subsequent training stage, but also leads to a significant reduction in baseband sampling overhead.
@@
However, it should be noted that accomplishing localization within a single sensing time slot generally necessitates extra baseband samples that are comparable in number to the size of the antenna array [23], which is extremely large in XL-MIMO systems. To address this problem, in this paper we propose to leverage a dual-functional XL-MIMO to promote efficient near-field CE. As shown in Fig. 1(b), radio frequency (RF) switches are adopted on antenna elements, which selectively activate either the $S$ branch for sensing or the $T$ branch for training. The communication part is a typical hybrid analog and digital architecture for XL-MIMO arrays [8], while the sensing part is composed of off-the-shelf components, including phase shifters, combiners, local oscillator, and power sensors, cf. Fig. 3.
Compared to the conventional MIMO transceiver architecture, the sensing component serves as the differentiating hardware. A similar architecture has been extensively utilized in holographic transceivers for microwave sensing [30] and imaging [31] within the near-field region, primarily due to its low-cost and plug-and-play characteristics. Specifically, in an effort to improve the integration efficiency and reduce the baseband sampling, power sensors [32] are considered. Instead of energy-consuming and costly high-frequency ADCs and baseband samplers, power sensors measure the power pattern of the received wave combined with the phase shifted reference wave in an energy-efficient and economical manner [33, 34]. An initial attempt to incorporate power sensor-based antenna elements into wireless communications was made in [29] for RIS-assisted systems, where efficient parameter estimation and beamforming algorithms were developed. In this paper, we shall also leverage this practical hardware architecture, based on which baseband sampling-free localization is designed.
@@ II-C Sensing Model
This subsection consider the first time slot when the RF switches equipped at the BS are set to the $S$ branch, such that the sensing function is activated. Based on the uplink sensing signal transmitted from the UE, the received object wave at the $n$-th antenna element of the BS is described as
@@
where ${\bf s}_{\rm S}\in\mathbb{C}^{N_{\rm UE}}$ denotes the sensing signal. Without loss of generality, we consider the case when ${\bf s}_{\rm S}$ is an all-one vector [24]. As depicted in Fig. 3, the received object waves are then combined with the reference wave as ${\bf e}_{\rm ref}[n]+{\bf e}_{\rm obj}[n]$, where ${\bf e}_{\rm ref}=A[e^{\jmath\xi_{1}},\cdots,e^{\jmath\xi_{N_{\rm BS}}}]^{T}$ is the phase shifted signal from local oscillator with $A$ denoting the manipulated amplitude of the reference wave and $\xi_{n}$ denoting the phase shift at the $n$-th antenna element of the BS array. Then, the $n$-th power sensor [29, 30] deployed at the XL-MIMO transceiver measures the power pattern of combined signals as
@@ II-D Pilot Training Model
Next, we consider the subsequent uplink pilot training stage, when the UE transmits the remaining pilot signals ${\bf s}_{\rm P}^{(t)}\in\mathbb{C}^{N_{\rm UE}^{\rm S}}$ ($1\!\leq\!t\!\leq\!\tau$) for CE. To reduce the number of the pricey and power-hungry RF chains in XL-MIMO systems, we consider hybrid transceiver architectures [13, 14, 15] for both the BS and UE. In this regard, the received signal at the BS in the $t$-th time slot is expressed as
@@
where ${\bf W}_{\rm RF}^{(t)}\in\mathbb{C}^{N_{\rm BS}\times N^{\rm RF}_{\rm BS}}$ and ${\bf W}_{\rm BB}^{(t)}\in\mathbb{C}^{N^{\rm RF}_{\rm BS}\times N^{\rm S}_{\rm BS}}$ denote the analog and digital combining matrices, and ${\bf F}_{\rm RF}^{(t)}\in\mathbb{C}^{N_{\rm UE}\times N^{\rm RF}_{\rm UE}}$ and ${\bf F}_{\rm BB}^{(t)}\in\mathbb{C}^{N^{\rm RF}_{\rm UE}\times N^{\rm S}_{\rm UE}}$ denote the analog and digital precoding matrices, respectively. Here, $N_{\rm UE}^{\rm RF}$ ($N_{\rm BS}^{\rm RF}$) and $N_{\rm UE}^{\rm S}$ ($N_{\rm BS}^{\rm S}$) denote the numbers of RF chains and data streams at the UE (BS), respectively, ${\bf n}^{(t)}\sim\mathcal{CN}(0,\sigma_{\rm n}^{2}{\bf I})$ is the additive white Gaussian noise (AWGN) vector, and $\sigma_{n}^{2}$ is the power of the AWGN at each receive antenna.
@@ IV-A Time Inversion Near-Field Reconstruction
Since the power of the reference signal in (7) contains no information about the UE and scatterers, we firstly subtract it and define the hologram ${\bm{\hbar}}\in\mathbb{R}^{N_{\rm BS}}$ [36] from the power measurement as
@@
where the object power pattern $\left|{\bf e}_{\rm obj}[n]\right|^{2}\ll\left|{\bf e}_{\rm ref}[n]\right|^{2}$ is neglected [^3] Note that the power of the reference wave can be artificially controlled and can be obtained precisely during the calibration procedure. On the other hand, the power of the object wave is inversely proportional to the square of distance between UE and BS arrays, and is far lower than that of the power of reference wave.. Note that the phase information of the object wave was discarded in (11). To address this issue, we multiply the hologram ${\bm{\hbar}}$ by the coherent reference wave as
@@
Then, we can emulate the time inversion EM wave propagation back to any given coordinate ${\bf r}=(x,y)$ by constructing the following function
@@
denote the desired term that contributes to the reconstruction and the interference term introduced by the the power sampling procedure (7) of the $\ell$-th path, respectively.
@@
[…]

### v3

@@ Abstract
Future sixth-generation (6G) systems are expected to leverage extremely large-scale multiple-input multiple-output (XL-MIMO) technology, which significantly expands the range of the near-field region. The spherical wavefront characteristics in the near field introduce additional degrees of freedom (DoFs), namely distance and angle, into the channel model, which leads to unique challenges in channel estimation (CE). In this paper, we propose a new sensing-enhanced uplink CE scheme for near-field XL-MIMO, which notably reduces the required quantity of baseband samples and the dictionary size. In particular, we first propose a sensing method that can be accomplished in a single time slot. It employs power sensors embedded within the antenna elements to measure the received power pattern rather than baseband samples. A time inversion algorithm is then proposed to precisely estimate the locations of users and scatterers, which offers a substantially lower computational complexity. Based on the estimated locations from sensing, a novel dictionary is then proposed by considering the eigen-problem based on the near-field transmission model, which facilitates efficient near-field CE with less baseband sampling and a more lightweight dictionary. Moreover, we derive the general form of the eigenvectors associated with the near-field channel matrix, revealing their noteworthy connection to the discrete prolate spheroidal sequence (DPSS). Simulation results unveil that the proposed time inversion algorithm achieves accurate localization with power measurements only, and remarkably outperforms various widely-adopted algorithms in terms of computational complexity. Furthermore, the proposed eigen-dictionary considerably improves the accuracy in CE with a compact dictionary size and a drastic reduction in baseband samples by up to 66%.
@@ I-A Related Works
Notably, an ideal dictionary needs to conform to the spatial characteristics of the channel in both distance and angle domains. This stipulation implies that acquiring localization information for both user equipments (UEs) and scatterers is a fundamental task in dictionary design for near-field CE [18, 19]. Specifically, when prior localization information pertaining to the propagation environment is incorporated into the dictionary, it can significantly reduce the need for further pilots and subsequent baseband samples for near-field CE. In recent years, integrated sensing and communication (ISAC) has made significant progress, posing sensing as a promising functionality in future communication systems [20]. Thanks to the large aperture provided by XL-MIMO arrays, efficient location estimation [21] as well as movement tracking [22] of UEs and scatterers become possible. This motivates us to customize a novel sensing-enhanced CE scheme for near-field XL-MIMO.
There have been some initial attempts on localization via XL-MIMO or sensing-assisted CE. Near-field radar sensing utilizing orthogonal frequency division multiplexing (OFDM) waveforms was validated in [21], while [22] proposed a near-field tracking method by measuring the curvature of arrival of the spherical wavefront from the UE. Besides, a localization method based on the multiple signal classification (MUSIC) algorithm has also been proposed to decouple the estimation of angles of arrival (AoAs) and distances of UEs located in the near field [23]. However, an excessive number of additional baseband samples are required for localization when directly integrating such methods into existing CE schemes, which counteracts the reduction of samples for CE brought by localization. Furthermore, a ping-pong pilot-based active-sensing method was proposed to achieve near-field beam alignment between the base station (BS) and the UE[24], where neural networks are utilized for beamforming optimization. Nevertheless, the pilot overhead and synchronization issues inherent in the alternating uplink and downlink training procedure remain challenging. To further improve the utilization of pilot signals, a sensing-assisted CE scheme was then proposed in [25]. While the pilot signals are reused for both localization and CE procedures and therefore no baseband sampling is required for localization, fully-digital XL-MIMO arrays are essentially the prerequisite of existing sensing-assisted CE. This array architecture requires a large number of baseband samples (typically dozens of times the number of antennas on the XL-MIMO array) via power-hungry analog-to-digital converters (ADCs) and, hence, is far from practical. Furthermore, the computational complexity of classical localization algorithms, e.g., MUSIC, applied in [23] and [25] is exceptionally high, which is caused by massive matrix multiplications for super-resolution grid search on both distance and angle domains. Therefore, pragmatic CE schemes enhanced by localization with smaller sampling overhead and lower computational complexity still require further exploration.
@@ II System Model
In this section, we firstly describe the considered near-field transmission scenario and present the near-field channel model. Furthermore, the uplink sensing-enhanced CE protocol is illustrated. The power sensor-based and hybrid antenna array is then introduced, whereas the corresponding uplink sensing and training models are further provided.
@@ II-A Channel Model
Consider a scenario with a UE array [^1] The proposed design in this paper can be extended to multi-user scenarios by assigning orthogonal pilots for different UEs. and a BS equipped with a hybrid analog-digital architecture XL-MIMO array operating in the TDD mode within its near-field region, where both the UE and BS are equipped with uniform linear arrays (ULAs) [^2] We consider ULA here for brevity, while it can be generalized to other antenna geometries. For example, it can be extended to uniform planar array (UPA) by applying Kronecker products to steering vectors in (2)., as shown in Fig. 1(a). The Cartesian coordinates of the $m$-th element of the UE’s antenna array and the $n$-th element of the BS antenna array are denoted by ${\bf r}_{\rm UE}^{(m)}=(x_{\rm UE}^{(m)},y_{\rm UE}^{(m)})^{T}$ and ${\bf r}_{\rm BS}^{(n)}=(x_{\rm BS}^{(n)},0)^{T}$, respectively. For uni-polarized antenna elements, the spherical wave spatial impulse response in the scalar form [27] is given by
@@ II-B Sensing-Enhanced CE Protocol and Power Senor-based XL-MIMO
We consider the uplink CE at the BS, which can be used to facilitate the downlink transmission. Conventionally, the UE transmits $\tau$ uplink pilot signals during the training stage, while the BS equipped with hybrid analog-digital architecture estimates the channel based on the received pilots [29, 13]. However, the pilot length $\tau$ and the resulting number of baseband samples at the BS can be exceedingly large due to limited RF chains in the hybrid array architecture and the complicated near-field channel model [18, 30].
In this paper, we propose a new sensing-enhanced CE protocol, where a single sensing time slot is inserted before the conventional pilot signal transmission, as shown in Fig. 2. In particular, in this additional slot, the RF switch is set to the $S$ branch (cf. Fig. 1(b)), enabling the UE to transmit an uplink sensing signal to realize the localization of both the UE and scatterers. In the subsequent $\tau$ slots, the RF switch is then set to the $T$ branch, where the UE transmits uplink pilots to facilitate CE. The main benefit of adding such a sensing time slot is that the estimated location coordinates can provide accurate prior localization information, thereby facilitating efficient near-field CE. This approach effectively reduces the pilot length $\tau$ in the subsequent training stage.
@@
However, it should be noted that conventional localization techniques generally necessitate extra baseband samples that are often several times the size of antenna arrays [25]. In fact, each baseband sample is generated through an RF front end with a high-precision ADC, whose power consumption increases exponentially with the number of quantization bits, and can reach up to the watt level in XL-MIMO systems [33]. This forms the first obstacle for practical localization methods in the baseband. Moreover, in the widely considered hybrid analog-digital array architecture [8], the number of RF chains and the corresponding ADCs is much fewer than that of antenna elements. Therefore, more time slots are needed to obtain a sufficient number of baseband measurements for accurate localization, which leaves significantly fewer time slots available for subsequent CE.
To address this problem, in this paper we propose a power sensor-based localization scheme in the RF domain. It can be accomplished in only one sensing slot and promote efficient near-field CE. As shown in Fig. 3, the sensing part, composed of off-the-shelf hardware components including phase shifters, combiners, local oscillator, and power sensors, distinguishing it from other setups [34]. Rather than relying on energy-consuming and costly high-frequency ADCs and baseband samplers, this approach utilizes power sensors to measure the power pattern of the received wave combined with the phase-shifted reference wave in an energy-efficient (down to milliwatt level) and economical manner [35, 36].
A similar architecture has been extensively utilized in holographic transceivers for microwave sensing [32, 37] and imaging [38] within the near-field region, primarily due to its low-cost and plug-and-play characteristics. Recently, an initial attempt to incorporate power sensor-based antenna elements into wireless communications was made in [31] for RIS-assisted systems, where efficient parameter estimation and beamforming algorithms were developed. In this paper, we shall also leverage this practical hardware architecture, based on which baseband sampling-free localization is designed.
@@ II-C Sensing Model
This subsection considers the first time slot when the RF switches equipped at the BS are set to the $S$ branch, such that the sensing function is activated. Based on the uplink sensing signal transmitted from the UE, the received object wave at the $n$-th antenna element of the BS is described as
@@
where ${\bf s}_{\rm S}\in\mathbb{C}^{N_{\rm UE}}$ denotes the sensing signal. Without loss of generality, we consider the case when ${\bf s}_{\rm S}$ is an all-one vector [26]. As depicted in Fig. 3, the received object waves are then combined with the reference wave as ${\bf e}_{\rm ref}[n]+{\bf e}_{\rm obj}[n]$, where ${\bf e}_{\rm ref}=A[e^{\jmath\xi_{1}},\cdots,e^{\jmath\xi_{N_{\rm BS}}}]^{T}$ is the phase-shifted signal from the local oscillator with $A$ denoting the manipulated amplitude of the reference wave and $\xi_{n}$ denoting the phase shift at the $n$-th antenna element of the BS array. Then, the $n$-th power sensor [31, 32] deployed at the XL-MIMO transceiver measures the power pattern of combined signals as
@@ II-D Pilot Training Model
Next, we consider the subsequent uplink pilot training stage, when the UE transmits the remaining pilot signals ${\bf s}_{\rm P}^{(t)}\in\mathbb{C}^{N_{\rm UE}^{\rm S}}$ ($1\!\leq\!t\!\leq\!\tau$) for CE. To reduce the number of pricey and power-hungry RF chains in XL-MIMO systems, we consider hybrid transceiver architectures [13, 14, 15] for both the BS and UE. In this regard, the received signal at the BS in the $t$-th time slot is expressed as
@@
where ${\bf W}_{\rm RF}^{(t)}\in\mathbb{C}^{N_{\rm BS}\times N^{\rm RF}_{\rm BS}}$ and ${\bf W}_{\rm BB}^{(t)}\in\mathbb{C}^{N^{\rm RF}_{\rm BS}\times N^{\rm S}_{\rm BS}}$ denote the analog and digital combining matrices, and ${\bf F}_{\rm RF}^{(t)}\in\mathbb{C}^{N_{\rm UE}\times N^{\rm RF}_{\rm UE}}$ and ${\bf F}_{\rm BB}^{(t)}\in\mathbb{C}^{N^{\rm RF}_{\rm UE}\times N^{\rm S}_{\rm UE}}$ denote the analog and digital precoding matrices, respectively. Here, $N_{\rm UE}^{\rm RF}$ ($N_{\rm BS}^{\rm RF}$) and $N_{\rm UE}^{\rm S}$ ($N_{\rm BS}^{\rm S}$) denote the numbers of RF chains and data streams at the UE (BS), respectively, ${\bf n}^{(t)}\sim\mathcal{CN}(0,\sigma_{\rm n}^{2}{\bf I})$ is the additive white Gaussian noise (AWGN) vector, and $\sigma_{\rm n}^{2}$ is the power of the AWGN at each receive antenna.
@@ IV-A Time Inversion Near-Field Reconstruction
Since the power of the reference signal ${\bf e}_{\rm ref}$ itself in (7) contains no information about the UE and scatterers, we firstly subtract it and define the hologram ${\bm{\hbar}}\in\mathbb{R}^{N_{\rm BS}}$ [40] from the power measurement as
@@
where the object power pattern $\left|{\bf e}_{\rm obj}[n]\right|^{2}\ll\left|{\bf e}_{\rm ref}[n]\right|^{2}$ is neglected [^3] Note that the power of the reference wave can be artificially controlled and can be obtained precisely during the calibration procedure. On the other hand, the power of the object wave is inversely proportional to the square of the distance between UE and BS arrays, and is far lower than that of the power of the reference wave.. Note that the phase information of the object wave was discarded in (11). To address this issue, we multiply the hologram ${\bm{\hbar}}$ by the coherent reference wave as
@@
Then, we can emulate the time inversion EM wave propagation back to any given coordinate [^4] It is a near-field generalization of the matched filter method, which degenerates into the matched filter method in the far-field region. ${\bf r}=(x,y)$ by constructing the following function
@@
denote the desired term that contributes to the reconstruction and the interference term introduced by the power sampling procedure (7) of the $\ell$-th path, respectively.
@@
$\lim_{\begin{subarray}{c}\kappa\to{\infty}\\ N_{\rm BS}\to{\infty}\end{subarray}}\!\!\left|\tilde{E}_{\ell}^{\rm d}({\bf r})\right|\!=\!\begin{cases}A^{2}N_{\rm BS}^{2}\sum_{n=1}^{N_{\rm BS}}\frac{1}{\left(d^{(\ell,n)}\right)^{2}}\!\!\!\!&\text{if}~{}{\bf r}={\bf r}_{\rm S}^{(\ell)}\!,\\ 0&\text{otherwise}.\end{cases}$ (22)
@@
Remark 2.
In XL-MIMO systems, spatial non-stationarity can occur, meaning that some multipath components are detected only by certain antennas, but not available for other antennas. Our proposed method is also applicable to these potential spatial non-stationary scenario, although the reduced effective aperture of the visible array may lead to higher localization errors. With the widely adopted orthogonal frequency division multiplexing (OFDM) waveform, we can aggregate the reconstruction results over multiple subcarriers, which effectively improves the distance resolution and suppresses the interference to relieve the performance drain caused by spatial non-stationarity [41].